# Univariate-Guided Sparse Regression for Biobank-Scale High-Dimensional -omics Data

### v1

@@ 4 Use of External Scores
The idea is to use the univariate coefficients from $\tt E$ rather than compute the LOO estimates from the training data. In the equation above these correspond to the $\tilde{\beta}_{j}$ terms. Specifically, the magnitudes $|\tilde{\beta}_{j}|$ serve as adaptive weights in the $\ell_{1}$-penalty term, while the signs of $\tilde{\beta}_{j}$ agree with the signs of the estimated coefficients $\gamma_{j}$ (or they are zero). We draw the univariate slopes and intercepts from the UK Biobank training data, but use the external coefficients $\tilde{\beta}_{j}$ to define the adaptive penalty weights and sign constraints in the second-stage regression.
The final model can be written as $\hat{\eta}(x)=\hat{\theta}_{0}+\sum_{j=1}^{p}\hat{\theta}_{j}x_{j}$, with $\hat{\theta}_{j}=\hat{\beta}_{j}\hat{\gamma}_{j}$, and $\hat{\theta}_{0}=\hat{\gamma}_{0}+\sum_{\ell=1}^{p}\hat{\beta}_{0\ell}\hat{\gamma}_{\ell}$. Here, we reincorporate the univariate slopes and intercepts from the first stage to obtain the final fitted model. The parallels with the standard uniLasso formulation are apparent. Note this alternative construction holds for binary responses as well.
@@
Table 3: Proportion of shared nonzero coefficients between each pair of models (Lasso (L), uniLasso (U), and uniLasso with external scores (E)) for all phenotypes. Each value represents the size of the intersection divided by the size of the union of non-zero SNPs.
@@ 5 Comparison of uniLasso to Other Approaches
In this section we summarize the performance differences among the methods considered. In addition to the three Lasso-based methods we analyze PRS-CS [Ge et al., 2019], a summary statistics-based Bayesian regression method. PRS-CS computes polygenic risk scores using GWAS summary statistics and an external LD reference panel, but it neither incorporates external summary statistics nor induces sparsity in its effect-size estimates. For PRS-CS, we first generated GWAS summary statistics on the combined training and validation (n=269,711)(n=269{,}711) with age, sex, and the top 10 PCs as covariates using PLINK v2.0a7 [Chang et al., 2015, Purcell and Chang, 2020]. We then applied PRS-CS with default settings using the precomputed European-ancestry LD reference panel derived from the 1000 genomes samples (PRScs GitHub Repo). Finally, we extracted the posterior effect-size estimates and computed polygenic risk scores using PLINK2’s --score subcommand [Chang et al., 2015].
PRS-CS underperforms the three Lasso-based methods on three of the four phenotypes and matches uniLasso ES on CHD (see Table 4). Combined with its higher computational cost and lack of sparsity, this limits its suitability for settings where interpretability is important. This makes it harder to gain any intuition behind which variants are actually driving the signal.
@@
BMI 0.119 0.103 0.118 0.070
Asthma 0.619 0.620 0.623 0.596
CHD 0.758 0.757 0.760 0.760
Table 4: Test set predictive performance of the four PRS methods for each phenotype. For Height and BMI, the metric is $R^{2}$; for Asthma and Coronary Heart Disease (CHD), the metric is AUC.
@@ 6 Discussion
We evaluate our methods on SNP data to demonstrate their gains in sparsity and interpretability over existing approaches. Our experiments indicate that uniLasso achieves about 40% more sparsity than Lasso while maintaining comparable predictive performance. Furthermore, integrating external data into the framework yields a noteworthy improvement in prediction accuracy. In the UK Biobank application, uniLasso with external scores (uniLasso ES) achieves the best or nearly best predictive performance across all four phenotypes. Additionally, uniLasso ES retains substantially fewer non-zero coefficients than Lasso and far fewer than PRS-CS, which keeps all SNPs in the model.
@@
Across all settings, uniLasso’s enforced sparsity enables efficient computation in ultra–high-dimensional regimes while preserving, and often improving upon, the predictive accuracy of dense alternatives. It is particularly striking that for height, a canonical highly polygenic trait, our best uniLasso models rely on only 30,00040,00030{,}000-40{,}000 non-zero SNPs out of more than one million candidates, yet still achieve test $R^{2}\approx 0.71$, clearly outperforming PRS-CS on the same dataset. Despite the extensive LD structure across the genome, uniLasso appears able to concentrate predictive signal onto a comparatively small subset of variants or LD blocks. This pattern is consistent with a view in which much of the SNP-heritability is effectively compressed into a modest number of genomic locations, even for traits that are often described as “highly polygenic”.
@@
From this perspective, our results provide an empirical counterpoint to strictly “pangenomic” or infinitesimal models in which every variant is assumed to have a non-zero effect. Dense approaches like LDpred-inf and PRS-CS remain valuable, especially when only GWAS summary statistics are available, but our findings suggest that (at least for traits like height) there is little predictive benefit in keeping millions of non-zero coefficients once LD and sample size are properly accounted for. A sparse-polygenic view, where a few tens of thousands of variants capture most of the predictive signal, appears sufficient for practical prediction in this setting. Methods such as PRS-CS or infinitesimal-prior approaches effectively assume that every variant has a non-zero effect, but that most effects are very small. In the presence of LD, these dense models spread each signal over many correlated SNPs, so that each individual predictor carries a heavily shrunk, “flattened” effect size. This is attractive for aggregate prediction and some forms of inference, but it makes it difficult to identify which specific variants or LD blocks are truly driving the score. Moreover, for individual-level risk prediction, relying solely on correlated tag SNPs has a causal cost: whenever the LD between tag and causal variant is imperfect, individuals whose genotypes deviate from the dominant LD pattern receive contributions from non-causal predictors that are essentially noise. In contrast, a sparse multivariate method that concentrates weight on (or near) truly causal variants aims to reduce this mismatch between the mechanistic signal and the predictors used for stratification, potentially improving both interpretability and robustness of individual risk estimates.
@@ Appendix A Algorithmic Runtimes
Method / Stage Continuous Response Binary Response
@@
   PRS-CS 5.5 hr 5.5 hr
Table 5: Approximate computation times for model training across all methods. Standard uniLasso includes an additional preprocessing stage in which the the LOO fitted values and univariate coefficients are computed for each feature.
@@
The PRS-CS pipeline involves substantial preprocessing and data formatting before the main model can be run, and additional steps are required afterward to aggregate the scores into a usable test metric. On Stanford’s Sherlock cluster, the primary PRS-CS Python command took roughly 5.5 hours to run, regardless of phenotype (Table 5). In our experiments it showed somewhat weaker predictive performance compared to the Lasso-based methods, and does not naturally yield sparse models.

### v4

@@ 4 Use of External Scores
The idea is to use the univariate coefficients from $\tt E$ rather than compute the LOO estimates from the training data. In the equation above these correspond to the $\tilde{\beta}_{j}$ terms. Specifically, the magnitudes $|\tilde{\beta}_{j}|$ serve as adaptive weights in the $\ell_{1}$-penalty term, while the signs of $\tilde{\beta}_{j}$ agree with the signs of the estimated coefficients $\gamma_{j}$ (or they are zero). We draw the univariate slopes and intercepts from the UK Biobank training data, but use the external coefficients $\tilde{\beta}_{j}$ to define the adaptive penalty weights and sign constraints in the second-stage regression. The final model can be written as $\hat{\eta}(x)=\hat{\gamma}_{0}+\sum_{j=1}^{p}\hat{\gamma}_{j}x_{j}.$ Note this alternative construction holds for binary responses as well.
@@
Table 3: Proportion of shared nonzero coefficients between each pair of models (Lasso (L), uniLasso (U), and uniLasso with external scores (E)) for all phenotypes. Each value represents the size of the intersection divided by the size of the union of non-zero SNPs. Recall that the binary phenotypes have much smaller active sets, as shown in Table 1.
@@ 5 Comparison of uniLasso to Other Approaches
In this section we summarize the performance differences among the methods considered. In addition to the three Lasso-based methods we analyze PRS-CS [Ge et al., 2019], a summary statistics-based Bayesian regression method. PRS-CS computes polygenic risk scores using GWAS summary statistics and an external LD reference panel, but it neither incorporates external summary statistics nor induces sparsity in its effect-size estimates. For PRS-CS, we first generated GWAS summary statistics on the combined training and validation sets (n=269,711)(n=269{,}711) with age, sex, and the top 10 PCs as covariates using PLINK v2.0a7 using MAF $\geq$ 0.0005 [Chang et al., 2015, Purcell and Chang, 2020]. We then applied PRS-CS with default settings using the precomputed European-ancestry LD reference panel derived from the 1000 genomes samples (PRScs GitHub Repo). Finally, we extracted the posterior effect-size estimates and computed polygenic risk scores using PLINK2’s --score subcommand [Chang et al., 2015].
PRS-CS underperforms the three Lasso-based methods on all four of the phenotypes, though it nearly matches uniLasso ES on CHD (see Table 4). Combined with its higher computational cost and lack of sparsity, this limits its suitability for settings where interpretability is important. In particular, the dense set of nonzero effects makes it harder to gain intuition behind which variants are actually driving the signal.
@@
BMI 0.119 0.103 0.118 0.071
Asthma 0.619 0.620 0.623 0.597
CHD 0.758 0.757 0.760 0.759
Table 4: Test set predictive performance of the four PRS methods for each phenotype. For Height and BMI, the metric is $R^{2}$; for Asthma and Coronary Heart Disease (CHD), the metric is AUC. uniLasso with external scores is denoted as uniLasso ES and PRS-CS is a competitive Bayesian regression method discussed in Section 5.
@@ 6 Discussion
We evaluate our methods on SNP data to demonstrate their gains in sparsity and interpretability over existing approaches. Our experiments indicate that uniLasso achieves about 40% more sparsity than Lasso while maintaining comparable predictive performance. Furthermore, integrating external data into the framework yields a noteworthy improvement in prediction accuracy. In the UK Biobank application, uniLasso with external scores (uniLasso ES) achieves the best predictive performance across all four measured phenotypes. Additionally, uniLasso ES retains substantially fewer non-zero coefficients than Lasso and far fewer than PRS-CS, which keeps all SNPs in the model.
@@
Across all settings, uniLasso’s enforced sparsity enables efficient computation in ultra–high-dimensional regimes while preserving, and often improving upon, the predictive accuracy of dense alternatives. It is particularly striking that for height, a canonical highly polygenic trait, our best uniLasso models rely on only about 34,00034{,}000 non-zero SNPs out of more than one million candidates, yet still achieve test $R^{2}\approx 0.71$. Despite the extensive LD structure across the genome, uniLasso appears able to concentrate predictive signal onto a comparatively small subset of variants or LD blocks. This pattern is consistent with a view in which much of the SNP-heritability is concentrated in a modest number of genomic locations, even for traits that are often described as “highly polygenic”.
@@
From this perspective, our results provide an empirical counterpoint to strictly “pangenomic” or infinitesimal models in which every variant is assumed to have a non-zero effect. Dense approaches like LDpred-inf and PRS-CS remain valuable, especially when only GWAS summary statistics are available, but our findings suggest that there is little predictive benefit in keeping millions of non-zero coefficients once LD and sample size are properly accounted for. A sparse-polygenic view, where a few tens of thousands of variants capture most of the predictive signal, appears sufficient for practical prediction in this setting. Methods such as PRS-CS or infinitesimal-prior approaches effectively assume that every variant has a non-zero effect, but that most effects are very small. In the presence of LD, these dense models spread each signal over many correlated SNPs, so that each individual predictor carries a heavily shrunk, “flattened” effect size. This is attractive for aggregate prediction and some forms of inference, but it makes it difficult to identify which specific variants or LD blocks are truly driving the score. Moreover, for individual-level risk prediction, relying solely on correlated tag SNPs has a causal cost: whenever the LD between tag and causal variant is imperfect, individuals whose genotypes deviate from the dominant LD pattern receive contributions from non-causal predictors that are essentially noise. In contrast, a sparse multivariate method that concentrates weight on (or near) truly causal variants aims to reduce this mismatch between the mechanistic signal and the predictors used for stratification, potentially improving both interpretability and robustness of individual risk estimates.
@@ Appendix A Algorithmic Runtimes
Method / Stage Cont. Response Bin. Response
@@
   PRS-CS 5.5 hr 2.5 hr
Table 5: Approximate computation times for model training across all methods for continuous and binary phenotypes. Standard uniLasso includes an additional preprocessing stage in which the the LOO fitted values and univariate coefficients are computed for each feature. PRS-CS also requires a GWAS as input, which we count as preprocessing.
@@
The PRS-CS pipeline involves substantial preprocessing and data formatting before the main model can be run, and additional steps are required afterward to aggregate the scores into a usable test metric. On Stanford’s Sherlock cluster, the primary PRS-CS Python command required roughly 5.5 hours for continuous phenotypes and 2.5 hours for binary phenotypes (Table 5). PRS-CS requires GWAS summary statistics as input, however, so we first run a genome-wide association study to generate the necessary single-variant effect estimates. This preprocessing stage showed the opposite pattern: about 1 hour for continuous responses versus 5 hours for binary. Overall, the end-to-end runtimes were similar, totaling approximately 6.5 hours and 7.5 hours for continuous and binary phenotypes respectively. In our experiments it showed somewhat weaker predictive performance compared to the Lasso-based methods, and does not naturally yield sparse models.